OpenStudy (megannicole51):

Does the series converge? series of (2n)!/(n!)^2 from n=1 to infinity

OpenStudy (megannicole51):

i got 1 so does that mean it diverges?

OpenStudy (megannicole51):

actually doesnt that mean i need to do another test bc getting 1 wont tell me....i did the ratio test!

OpenStudy (megannicole51):

@agent0smith

OpenStudy (agent0smith):

1 means it's inconclusive, doesn't tell you anything. Ugh...this looks familiar but idk what test to do...

OpenStudy (megannicole51):

ill look in my book and see what i can find!

OpenStudy (agent0smith):

Wolfram alpha says try the limit test, it doesn't converge

OpenStudy (megannicole51):

what would my bn be for the limit test?

OpenStudy (megannicole51):

would it be 2n/n?

OpenStudy (agent0smith):

Not sure :/

OpenStudy (anonymous):

i don't think the terms go to zero

OpenStudy (megannicole51):

i figured it out:) thanks guys! i have one with a square root that I'm stuck on ill repost!! help if you can:)

OpenStudy (anonymous):

try it with a few numbers, small ones like \(5\) you would get \(\frac{10!}{(5!)^2}=\frac{10\times 9\times 8\times 7\times 6}{5\times 4\times 3\times 2}=\binom{10}{5}\)

OpenStudy (megannicole51):

yeah thats what i did!:) thanks!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

in fact, i am almost certain that \[\frac{(2n)!}{(n!)^2}=\binom{2n}{n}\]